OpenStudy (anonymous):

Find -3a2 - b3 + 3c2 - 2b3 if a = 2, b = -1, and c = 3.

OpenStudy (anonymous):

can anyone help me

OpenStudy (aripotta):

just plug them in -3(2)^2 - (-3)^3 + 3(3)^2 - 2(-1)^3

OpenStudy (aripotta):

-6^2 + 3^3 + 9^2 + 2^3

OpenStudy (aripotta):

36 + 27 + 81 + 8 = 152

OpenStudy (anonymous):

it is right

OpenStudy (anonymous):

x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3.

OpenStudy (anonymous):

x - 3y if x = 3 and y = -2. i belive this one is -9

OpenStudy (confluxepic):

That would be 3-(-6)=9.

OpenStudy (anonymous):

thank you can you help me with this one x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3.

OpenStudy (anonymous):

@Jesstho.-.

OpenStudy (anonymous):

Close this and open another one, you can't post two questions in the same thread @myininaya

OpenStudy (anonymous):

Ayee will you help me with the question

OpenStudy (anonymous):

yeah, close this then ask another question, tag me.

OpenStudy (anonymous):

x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3. @Jesstho.-.